How much energy (in kilojoules) is needed to heat 5.40 g of ice from -14.0 ∘C to 31.0 ∘C? The heat of fusion of water is 6.01kJ/mol, and the molar heat capacity is 36.6 J/(K⋅mol) for ice and 75.3 J/(K⋅mol) for liquid water.
1
Expert's answer
2019-09-09T04:59:46-0400

Number of moles of "5.4 gm" ice"=\frac{5.4}{18}=0.3"

Heat required to raise temperature from "-14\degree" to zero"=nc \delta t=0.3 \times" "36.6 \times 14 \times 10^{-3}=153.72 \times 10^{-3} J"

Heat required for fusion"=6.01 \times 10^{3} \times 0.3=1.803 \times 10^3J"

Heat required to raise temperature from zero to "31 \degree C=31 \times 75.3 \times 10^{-3} \times 0.3=700.29 \times 10^{-3}"

Total heat required"=1.803 \times 10^3+153.72 \times 10^{-3}+700.29 \times 10^{-3}=" "1803+0.854=1803.854 J\approx 1804 J"